OpenStudy (maddy1251):

A surveyor measures the angle of elevation at 20 degrees for a point 5 miles away. What is the vertical change in elevation from the point where the surveyor is standing to the point 5 miles away?

OpenStudy (flcon04):

What do you think it is?

OpenStudy (maddy1251):

4Is this somewhat right?

OpenStudy (flcon04):

Yeah.

OpenStudy (flcon04):

Ok give me a sec i will explain it step by step.

OpenStudy (flcon04):

That is how it should look like.

OpenStudy (flcon04):

from the triangle you do tan20 =y/5

OpenStudy (flcon04):

now what do you think the answer is?

OpenStudy (maddy1251):

tan20=y/5 0.36 = y/5 1.8=y @flcon04

OpenStudy (flcon04):

Correct!